OpenStudy (anonymous):

explain, in complete sentences, which method would you use to solve the following system of equations. Also explain why this method was chosen. provide solution. x-3y+4z=-9 2x+y+3z=6 x-4y+3z=-9

jhonyy9 (jhonyy9):

subtract 3rd from first equation x-3y+4z=-9 x-4y+3z=-9 ------------ 0 y+z=0

jhonyy9 (jhonyy9):

like first step

jhonyy9 (jhonyy9):

like second step multiply first by -2 and assum with second -2x+6y-8z=18 2x+y+3z=6 --------------- 0 7y-5z=24

jhonyy9 (jhonyy9):

so than we know tha y+z=0 so y=-z so this substitute in 7y-5z=24 and will get 7(-z)-5z=24 so -7z-5z=24 so -12z=24 so z=24/(-12) = -2 so z=-2 so y=-(-2)=2 2x+y+3z=6 2x+2+3(-2)=6 2x+2-6=6 2x-4=6 2x=10 x=10/2 x=5 x=5 y=2 z=-2